OpenStudy (anonymous):

Determine which of the lines, if any, are parallel or perpendicular. Line a passes through (-2,1) and (0,3) Line b passes through (4,1) and (6,4) Line c passes through (1,3) and (4,1)

OpenStudy (anonymous):

i think a and c are parallel

OpenStudy (aripotta):

you know the slope formula, right?\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] find the slope for each line first

OpenStudy (aripotta):

can you tell me what you get?

OpenStudy (anonymous):

If I get the answer

OpenStudy (aripotta):

well i'm not gonna give you the answer if that's what you're asking? find the slope for each line and you can see which lines are parallel or perpendicular

OpenStudy (anonymous):

oh no, I know. I'm not asking for the answer im just saying that if i figure out the answer then ill tell you

OpenStudy (aripotta):

oh ok, sorry i misunderstood :) parallel lines have the same slope and perpendicular lines have slopes that are negative reciprocals

OpenStudy (aripotta):

so, for example, perpendicular lines would have slopes like 4 and -1/4

OpenStudy (anonymous):

oh ok

OpenStudy (aripotta):

so have you found out the slopes for each line?

OpenStudy (aripotta):

or still workin on it?

OpenStudy (anonymous):

I'm still confused on what im supposed to do it. Im VERY

OpenStudy (anonymous):

bad at math.

OpenStudy (aripotta):

ok this is the slope formula: \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] it's the difference of the y's divided by the difference of the x's

OpenStudy (aripotta):

i'll do line a for you

OpenStudy (aripotta):

line a has points (-2,1) and (0,3). so i'm going to plug that into the formula\[\frac{ 1-3 }{ -2-0 }=\frac{ -2 }{ -2 }=1\] so line a has a slope of 1

OpenStudy (aripotta):

can you do line b?

OpenStudy (anonymous):

ok